# Continual Vision-and-Language Navigation

### v2

@@ 4.1 Perplexity Replay
To construct replay memory for continual learning, random sampling [44, 35] is often employed. However, this approach fails to consider the importance of samples in recovering past knowledge, resulting in suboptimal replay memory composition. Existing methods [39, 2] have attempted to address this by utilizing gradient-based or optimizable approaches to construct more effective replay memory. Nevertheless, these methods are typically tailored to simple tasks such as image classification and are not suitable for tasks that require sequential action prediction, such as CVLN.
Algorithm 1 Perplexity Replay Memory Update
1:  Input: Replay memory $M$, replay memory size $\lvert M\rvert$, set of episodes from the last scene domain $E_{\text{last}}$, agent parameters $\theta$, the number of trained scene domains $s$
2:  Output: Updated replay memory $M$
3:  Calculate Action Perplexity (AP) for each episode in $E_{\text{last}}$ and $M$
[…]
6:  $n_{s}\leftarrow\lvert M\rvert/(s+1)$
7:  for $s=1$ to $s$ do
8:     Remove episodes with the lowest AP from $M$ for domain $s$ until only $n_{s}$ episodes remain
9:  end for
10:  Add episodes from $E_{\text{last}}$ to $M$ until $\lvert M\rvert$ equals fixed size
11:  return  $M$
To overcome these limitations, we propose Perplexity Replay (PerpR), which constructs the replay memory by quantifying the difficulty of VLN episodes. A key innovation of PerpR is that the model self-evaluates the difficulty of each episode using Action Perplexity (AP), without relying on external metrics or annotations. By self-assessing its own uncertainty through AP, the model identifies challenging episodes where it is less confident. PerpR then prioritizes these difficult episodes in the replay memory, enabling the agent to focus on experiences that require further learning and thus enhancing the overall learning process.
@@
$\mathcal{L}_{D_{s_{c}}}$ $\triangleq\mathbb{E}_{(I,A^{*})\sim D_{s_{c}}}[\ell(\pi_{\theta}(V,I),A^{*})]$
$\mathcal{L}_{M}$ $\triangleq\mathbb{E}_{(I,A^{*})\sim M}[\ell(\pi_{\theta}(V,I),A^{*})]$
where $\mathcal{L}_{D_{s_{c}}}+\mathcal{L}_{M}$ represents the sum of the losses over the current scene domain $D_{s_{c}}$ and the replay memory $M$. The memory update procedure for PerpR is described in Algorithm 1.
@@ 4.2 Episodic Self-Replay
To effectively utilize the samples in the replay memory, prior works [34, 8, 7, 43] have employed logit distillation methods, which were designed for continual learning in non-sequential tasks such as image classification or object detection. However, they are not directly applicable to CVLN due to its sequential action prediction in dynamic environments. In CVLN, agents must perform sequences of actions where observations depend on previous actions, involving temporal state changes and decision-making processes. Existing logit distillation methods focus on single-step predictions, limiting their effectiveness in learning or generalizing complex sequential action patterns.
To address this limitation, we extended these approaches for application to CVLN episodes. At each action step, we stored the action logits in a replay memory. When learning from episodes in the replay memory, the CVLN agent utilizes these stored logits to train itself to follow its past action distribution at each step. To reinforce consistency, we applied student forcing [38] during training by using the stored logits as the next action for subsequent steps, rather than the predicted logits. This encourages adherence to prior action patterns and aligns the learning process with the sequential nature of CVLN.
Algorithm 2 Episodic Self-Replay
0:  Dataset $D_{s}$, Memory $M$, Learning rate $\eta$, Agent $f_{\theta}$, $\text{GetObservation}(t)$, $\text{ActionExecuter}(a_{t})$, $\lambda_{1}$, $\lambda_{2}$
1:  Sample $(I,A^{*})\sim D_{s}$
2:  $\mathcal{L}_{D_{c}}\leftarrow 0$, $Z\leftarrow[]$, $t\leftarrow 1$
3:  $v_{t}\leftarrow\text{GetObservation}(t)$;
4:  while $a_{t}$ does not predict stop do
5:     $z_{t}\leftarrow f_{\theta}(I,v_{t})$; $Z.\text{append}(z_{t})$
6:     $a_{t}\leftarrow\arg\max(z_{t})$
7:     $a_{t}^{*}\leftarrow A^{*}(t)$ //Get true action
8:     $\mathcal{L}_{D_{c}}\leftarrow\mathcal{L}_{D_{c}}+\ell(a_{t},a_{t}^{*})$
9:     $v_{t+1}\leftarrow\text{ActionExecuter}(a_{t})$
10:     $t\leftarrow t+1$
11:  end while
12:  $\mathcal{L}_{D_{c}}\leftarrow\mathcal{L}_{D_{c}}/t$
13:  Update $M\leftarrow\text{Random}(M,(I,A^{*},Z))$
14:  Sample $(I^{\prime},A^{\prime*},Z^{\prime*})\sim M$
15:  $\mathcal{L}_{ESR}\leftarrow 0$, $\mathcal{L}_{M}\leftarrow 0$, $t\leftarrow 1$
16:  $v_{t}\leftarrow\text{GetObservation}(t)$;
17:  while $a_{t}$ does not predict stop do
18:     $z_{t}\leftarrow f_{\theta}(I^{\prime},v_{t})$
19:     $a_{t}\leftarrow\arg\max(z_{t})$
20:     $a_{t}^{\prime*}\leftarrow A^{\prime*}(t)$ //Get true action
21:     $\mathcal{L}_{M}\leftarrow\mathcal{L}_{M}+\ell(a_{t},a_{t}^{\prime*})$
22:     $z_{t}^{\prime*}\leftarrow Z^{\prime*}(t)$ //Get replay logit
23:     $\mathcal{L}_{ESR}\leftarrow\mathcal{L}_{ESR}+\|z_{t}-z_{t}^{\prime*}\|^{2}_{F}$//Logit distillation
24:     $a_{t}^{*}\leftarrow\arg\max(z_{t}^{\prime*})$ //Get replay logit’s action
25:     $v_{t+1}\leftarrow\text{ActionExecuter}(a_{t}^{*})$
26:     $t\leftarrow t+1$
27:  end while
28:  $\mathcal{L}_{ESR}\leftarrow\mathcal{L}_{ESR}/t$
29:  $\mathcal{L}_{M}\leftarrow\mathcal{L}_{M}/t$
30:  $\mathcal{L}_{ESR\_{total}}\leftarrow\mathcal{L}_{D_{c}}+\lambda_{1}\mathcal{L}_{ESR}+\lambda_{2}\mathcal{L}_{M}$
31:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{ESR\_{total}}$
We define the agent parameters immediately after learning the scene domain $s$ as $\theta_{s}^{*}$. A function $f$ outputs the logits for action candidates at each sequential step of navigation. Unlike other rehearsal-based methods, ESR stores the sequential action logits $Z\triangleq\{z_{n}\}_{n=1}^{N}\triangleq f_{\theta_{s}^{*}}(V,I)$, corresponding to each step in the sequence, when updating the replay memory $M$. ESR’s replay memory uses random sampling [44] to select episodes to add and delete.
@@
$\mathcal{L}_{ESR}\triangleq\mathbb{E}_{(I,Z)\sim M}\left[\sum_{n=1}^{N}\|z_{n}-f_{\theta}(V_{n},I_{n})\|^{2}_{F}\right]$ (6)
Method I-CVLN D-CVLN
@@
1. Vanilla $23.1\pm 0.8$ $25.6\pm 0.6$ $10.5\pm 0.5$ $5.5\pm 0.4$
@@
where $\lambda_{1}$ and $\lambda_{2}$ control the impact of $\mathcal{L}_{M}$ and $\mathcal{L}_{ESR}$ on the overall loss. Both $\lambda_{1}$ and $\lambda_{2}$ are set to 0.2. Please refer to Algorithm 2 for more details.
5 Experiments
5.1 Comparative Methods
Random Replay (RandR) uses Reservoir sampling [44] to manage replay memory, integrating selected episodes with current training scene domain $D_{s_{c}}$.
A-GEM limits gradient updates to minimize memory loss, improving computational complexity over GEM.
AdapterCL adds domain-specific adapters [18]; the adapter with the lowest action perplexity is selected during evaluation.
Method I-CVLN D-CVLN
[…]
200 23.8 26.3 10.2 50 4.9
PerpR 500 23.8 26.4 9.8 100 6.2
1000 28.7 31.1 8.7 200 6.0
200 25.8 29.2 9.2 50 5.3
ESR 500 27.7 31.1 8.8 100 5.5
1000 29.8 34.4 9.1 200 5.9
Table 3: Comparative analysis of the impact of replay memory size on the performance for rehearsal-based methods. The table shows that larger memory sizes improve performance, indicating a positive correlation between buffer size and agent performance.
5.2 Experiment Setups
For I-CVLN, we use VLN-BERT [17] with Oscar [26] as the backbone, applying CL methods to VLN training. The rehearsal-based approach uses a fixed replay memory of 500, about 3% of the dataset, equally divided among scene domains. AdapterCL employs 20 adapters, matching the scene domains in I-CVLN, and adapters can increase with more domains. Three curricula with randomized scene domain orders are generated using distinct seeds, and results are averaged across them. For non-comparative analyses, results from a single curriculum are reported. For D-CVLN, we use a randomly initialized HAMT [11] as the backbone, applying CL methods to its VLN training. The replay memory size is set to 100, about 3% of the dataset, with 11 adapters fixed for AdapterCL. Three distinct curricula, each covering 11 scene domains, are configured, and results are averaged across these curricula.
@@ 5.3 Comparative Results
Performance Bounds: Vanilla and Joint provide performance bounds, where Vanilla reflects a lower bound due to a lack of mechanisms to mitigate forgetting, and Joint represents an upper bound with full access to all data simultaneously. The large performance gap between these two methods underscores the difficulty of applying continual learning to CVLN, where each new scene domain introduces unique challenges that increase the risk of forgetting prior knowledge.
@@
Our proposed methods provide a significant improvement in mitigating catastrophic forgetting and adapting to new environments, making them more suitable for complex, multi-domain CVLN tasks than traditional CL methods.
@@ 5.4 Analysis
5.4.1 Memory Size Analysis
The proposed methods are rehearsal-based methods that use replay memory. Since the performance of these methods can vary depending on the size of the replay memory, we examined the performance differences of rehearsal-based methods according to the size of their replay memory in Table 3. PerpR and ESR perform better as the replay memory size increases on both datasets. This demonstrates that larger memory can help retain previously learned knowledge by storing a wider variety of data samples.
5.4.2 Stability-Plasticity Trade-off Analysis
In CVLN, the agents need to have memory stability and learning plasticity. Memory stability ensures retention of previously acquired knowledge, while learning plasticity enables the efficient assimilation of new information. These attributes are often in trade-off, known as the stability-plasticity dilemma [33]. Stability (S) is defined as the agents’ average performance on previously encountered scene domains after training on a new one, while Plasticity (P) represents the average initial performance in newly encountered domains. Catastrophic forgetting occurs when plasticity exceeds stability. We evaluate the stability-plasticity trade-off [37], the harmonic mean of S and P, visualized in Figure 5. ESR achieves a better trade-off in I-CVLN than RandR and Vanilla, with greater plasticity and stability, while PerpR performs similarly in D-CVLN.
Ablation I-CVLN D-CVLN
@@ 5.4.2 Stability-Plasticity Trade-off Analysis
PerpR $\textbf{26.1}\pm 1.6$ $\textbf{28.9}\pm 1.8$ $\textbf{9.8}\pm 0.2$ $\textbf{6.1}\pm 0.4$
PerpR-Rev. $24.6\pm 1.1$ $27.1\pm 1.2$ $10.2\pm 0.3$ $5.4\pm 0.2$
Table 4: Impact of memory update reversal on performance in PerpR.
@@ 5.5 Ablation Study
5.5.1 Reversing PerpR memory update
Table 4 shows the results of reversing the replay memory update in PerpR. PerpR is developed based on the assumption that episodes with higher action perplexity would be more beneficial for future learning. The results confirm the validity of PerpR’s underlying assumption. PerpR’s performance degrades when replay memory updates are reversed by storing low action perplexity episodes and deleting high ones.
5.5.2 ESR loss ablation
Ablation I-CVLN D-CVLN
@@ 5.5.2 ESR loss ablation
$\mathcal{L}_{ESR\_total}$ 27.7 31.1 8.8 5.5
@@
Table 5: Ablation study on ESR for I-CVLN and D-CVLN. This table shows the effects of removing specific loss components on the overall performance.
Table 5 shows the results of an ablation study on the ESR loss in Equation 7. The removal of each of the two losses shows decreased performance across all datasets. Notably, the removal of $\mathcal{L}_{ESR}$ resulted in a greater performance drop than the removal of $\mathcal{L}_{M}$. This confirms that leveraging the action distribution predictions from previous models helps mitigate the catastrophic forgetting problem in CVLN agents. Additionally, we observe a synergistic effect when using both types of losses obtained through replay memory.
@@ 6 Conclusions and Future Work
In this paper, we introduced the Continual Vision-and-Language Navigation (CVLN) paradigm to enable VLN agents to learn in changing environments without forgetting previous knowledge. We proposed two settings, I-CVLN and D-CVLN, to account for various forms of natural language instruction. Additionally, we present two simple yet effective methods—Perplexity Replay and Episodic Self-Replay—that use rehearsal mechanisms to mitigate catastrophic forgetting and are applicable to tasks like CVLN, which involve sequential decision-making. Our experiments showed that these methods outperform existing continual learning approaches, highlighting the importance of effective replay memory in CVLN. For future work, we plan to extend CVLN to continuous environments and design new settings and datasets specifically for VLN continual learning, beyond existing dataset-based benchmarks.

### v3

@@ 4.1 Perplexity Replay
Random sampling [Vitter(1985), Riemer et al.(2018)Riemer, Cases, Ajemian, Liu, Rish, Tu, and Tesauro] is often used to construct replay memory for continual learning but overlooks sample importance for recovering past knowledge, leading to suboptimal memory composition. Existing methods [Shim et al.(2021)Shim, Mai, Jeong, Sanner, Kim, and Jang, Aljundi et al.(2019b)Aljundi, Lin, Goujaud, and Bengio] address this using gradient-based or optimizable approaches, but they are mainly designed for simple tasks like image classification and are unsuitable for sequential action prediction required in CVLN.
To address these limitations, we propose Perplexity Replay (PerpR), which builds replay memory by quantifying episode difficulty. PerpR introduces Action Perplexity (AP) to let the model self-evaluate episode difficulty without external metrics or annotations. Using AP, the model identifies challenging episodes with high uncertainty and prioritizes them in replay memory, focusing learning on harder experiences and improving overall performance.
@@
[…]
where $\mathcal{L}_{D_{s_{c}}}+\mathcal{L}_{M}$ represents the sum of the losses over the current scene domain $D_{s_{c}}$ and the replay memory $M$. The memory update procedure for PerpR is provided in the supplementary material (Section 4.1).
@@ 4.2 Episodic Self-Replay
To leverage replay memory effectively, prior works [Rebuffi et al.(2017)Rebuffi, Kolesnikov, Sperl, and Lampert, Buzzega et al.(2020)Buzzega, Boschini, Porrello, Abati, and Calderara, Boschini et al.(2022)Boschini, Bonicelli, Buzzega, Porrello, and Calderara, Verwimp et al.(2022)Verwimp, Yang, Parisot, Lanqing, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars] used logit distillation for non-sequential tasks like image classification. However, these methods are unsuitable for CVLN, which requires sequential action prediction. CVLN agents must handle temporally dependent observations and decision-making, while existing distillation methods focus only on single-step predictions, limiting their ability to learn complex action sequences.
To address this limitation, we extended these approaches for application to CVLN episodes. At each action step, we stored the action logits in a replay memory. When learning from episodes in the replay memory, the CVLN agent utilizes these stored logits to train itself to follow its past action distribution at each step. To reinforce consistency, we applied student forcing [Sauder et al.(2020)Sauder, Hu, Che, Mordido, Yang, and Meinel] during training by using the stored logits as the next action for subsequent steps, rather than the predicted logits. This encourages adherence to prior action patterns and aligns the learning process with the sequential nature of CVLN.
[…]
@@
where $\lambda_{1}$ and $\lambda_{2}$ control the impact of $\mathcal{L}_{M}$ and $\mathcal{L}_{ESR}$ on the overall loss. Both $\lambda_{1}$ and $\lambda_{2}$ are set to 0.2. Please refer to supplementary material (Section 4.2) for more details.
5 Experiments
5.1 Comparative Methods
Random Replay (RandR) uses Reservoir sampling [Vitter(1985)] to manage replay memory, integrating selected episodes with current training scene domain $D_{s_{c}}$.
A-GEM [Chaudhry et al.(2018)Chaudhry, Ranzato, Rohrbach, and Elhoseiny] limits gradient updates to minimize memory loss, improving computational complexity over GEM [Lopez-Paz and Ranzato(2017)].
AdapterCL adds domain-specific adapters [Houlsby et al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De Laroussilhe, Gesmundo, Attariyan, and Gelly]; the adapter with the lowest action perplexity is selected during evaluation.
5.2 Experiment Setups
For I-CVLN, we use VLN-BERT [Hong et al.(2020)Hong, Wu, Qi, Rodriguez-Opazo, and Gould] with Oscar [Li et al.(2020)Li, Yin, Li, Zhang, Hu, Zhang, Wang, Hu, Dong, Wei, et al.] as the backbone and apply CL methods during training. The rehearsal-based method uses a fixed replay memory of 500 samples (about 3% of the dataset), equally divided across 20 scene domains. AdapterCL employs 20 adapters matching these domains, expandable for more domains. Three curricula with different scene orders are generated using distinct seeds, and results are averaged; for non-comparative analyses, a single curriculum is used. For D-CVLN, we adopt a randomly initialized HAMT [Chen et al.(2021)Chen, Guhur, Schmid, and Laptev] backbone with CL methods. The replay memory is set to 100 samples (about 3% of the dataset), and AdapterCL uses 11 fixed adapters. Three curricula covering 11 scene domains are configured, and results are averaged.
[…]
@@ 5.3 Comparative Results
Performance Bounds: Vanilla and Joint define performance bounds, with Vanilla as the lower bound due to no forgetting mitigation, and Joint as the upper bound with full data access. Their large performance gap highlights the challenge of continual learning in CVLN, where new scene domains introduce unique challenges and increase forgetting risk.
@@ 5.4 Analysis
Memory Size Analysis The proposed rehearsal-based methods use replay memory, and their performance depends on its size. Table 2 shows that PerpR and ESR perform better with larger memory on both datasets, indicating that larger memory helps retain knowledge by storing more diverse samples.
@@
[…]
PerpR 500 23.8 26.4 9.8 100 6.2
1000 28.7 31.1 8.7 200 6.0
[…]
ESR 500 27.7 31.1 8.8 100 5.5
1000 29.8 34.4 9.1 200 5.9
[…]
Table 2: Comparative analysis of the impact of replay memory size on performance for rehearsal-based methods. Larger memory sizes improve performance, indicating a positive correlation between buffer size and agent performance.
Stability-Plasticity Trade-off Analysis In CVLN, agents require both memory stability to retain past knowledge and learning plasticity to acquire new information. These are often in trade-off, known as the stability-plasticity dilemma [Parisi et al.(2019)Parisi, Kemker, Part, Kanan, and Wermter]. Stability (S) measures average performance on past domains after learning a new one, while Plasticity (P) measures initial performance on new domains. Catastrophic forgetting occurs when plasticity outweighs stability. We assess this trade-off [Sarfraz et al.(2022)Sarfraz, Arani, and Zonooz] using the harmonic mean of S and P (Figure 4). ESR achieves a better trade-off in I-CVLN than RandR and Vanilla, while PerpR shows comparable performance in D-CVLN.
@@ 5.5 Ablation Study
Reversing PerpR memory update Table 4 shows the results of reversing the replay memory update in PerpR. PerpR is developed based on the assumption that episodes with higher action perplexity would be more beneficial for future learning. The results confirm the validity of PerpR’s underlying assumption. PerpR’s performance degrades when replay memory updates are reversed by storing low action perplexity episodes and deleting high ones.
@@
[…]
PerpR-Rev. $24.6\pm 1.1$ $27.1\pm 1.2$ $10.2\pm 0.3$ $5.4\pm 0.2$
[…]
Table 3: Impact of memory update reversal on performance in PerpR.
[…]
@@
[…]
Table 4: Ablation study on ESR for I-CVLN and D-CVLN. Effects of removing specific loss components.
ESR loss ablation Table 4 shows the results of an ablation study on the ESR loss in Equation 7. The removal of each of the two losses shows decreased performance across all datasets. Notably, the removal of $\mathcal{L}_{ESR}$ resulted in a greater performance drop than the removal of $\mathcal{L}_{M}$. This confirms that leveraging the action distribution predictions from previous models helps mitigate the catastrophic forgetting problem in CVLN agents. Additionally, we observe a synergistic effect when using both types of losses obtained through replay memory.
@@ 6 Conclusions and Future Work
In this paper, we introduced the Continual Vision-and-Language Navigation paradigm to enable VLN agents to learn in changing environments without forgetting previous knowledge. We proposed two settings, I-CVLN and D-CVLN, to account for various forms of natural language instruction. Additionally, we present two simple yet effective methods—Perplexity Replay and Episodic Self-Replay—that use rehearsal mechanisms to mitigate catastrophic forgetting and are applicable to tasks like CVLN, which involve sequential decision-making. Our experiments showed that these methods outperform existing continual learning approaches, highlighting the importance of effective replay memory in CVLN. For future work, we plan to extend CVLN to continuous environments and design new settings and datasets specifically for VLN, beyond existing dataset-based benchmarks.
Acknowledgements This work was partly supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) under grants RS-2021-II212068-AIHub (10%), RS-2021-II211343-GSAI (10%), RS-2022-II220951-LBA (15%), and RS-2022-II220953-PICA (15%); the National Research Foundation of Korea (NRF) under grants RS-2024-00353991-SPARC (15%) and RS-2023-00274280-HEI (15%); the Korea Evaluation Institute of Industrial Technology (KEIT) under grant RS-2024-00423940 (10%); and the Gwangju Metropolitan City (Artificial Intelligence Industrial Convergence Cluster Development Project, 10%) funded by the Korean government.